OpenStudy (anonymous):

Find the value of cos(7pi/12) using a half angle formula?

zepdrix (zepdrix):

Half angle formula for cosine: \[\huge \cos\left(\frac{x}{2}\right)=\sqrt{\frac{1+cosx}{2}}\] Let's try to put our angle in this form. \[\huge \cos \left(\frac{7\pi}{12}\right)=\cos \left(\frac{ (\frac{7\pi}{6}) }{ 2 }\right)\]

zepdrix (zepdrix):

\[\huge =\sqrt{\frac{1+\cos\left(\frac{7\pi}{6}\right)}{2}}\]

zepdrix (zepdrix):

From there, you just need to remember what the cos of 7pi/6 is, and simplify things down. I forget whether or not you need to worry about the +/- of the square root.. 7pi/12 is in the second quadrant, so I suppose we want the negative root.